# A Flexible Plug-and-Play Module for Generating Variable-Length

### v1

@@
Abstract.
Deep supervised hashing has become a pivotal technique in large-scale image retrieval, offering significant benefits in terms of storage and search efficiency. However, existing deep supervised hashing models predominantly focus on generating fixed-length hash codes. This approach fails to address the inherent trade-off between efficiency and effectiveness when using hash codes of varying lengths. To determine the optimal hash code length for a specific task, multiple models must be trained for different lengths, leading to increased training time and computational overhead. Furthermore, the current paradigm overlooks the potential relationships between hash codes of different lengths, limiting the overall effectiveness of the models. To address these challenges, we propose the Nested Hash Layer (NHL), a plug-and-play module designed for existing deep supervised hashing models. The NHL framework introduces a novel mechanism to simultaneously generate hash codes of varying lengths in a nested manner. To tackle the optimization conflicts arising from the multiple learning objectives associated with different code lengths, we further propose an adaptive weights strategy that dynamically monitors and adjusts gradients during training. Additionally, recognizing that the structural information in longer hash codes can provide valuable guidance for shorter hash codes, we develop a long-short cascade self-distillation method within the NHL to enhance the overall quality of the generated hash codes. Extensive experiments demonstrate that NHL not only accelerates the training process but also achieves superior retrieval performance across various deep hashing models. Our code is publicly available at https://github.com/hly1998/NHL.
Deep Hashing, Image Retrieval, Multiple Code Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
@@
With the accumulation of visual data on the Internet, the existing databases are becoming increasingly vast. To address the growth of data volume in large-scale image databases, hashing represents images as binary hash code for storage and search efficiency (Luo et al., 2023). Recently, deep supervised hashing has achieved significant advancements due to its ability to extract deep features from data and utilize supervised signals to ensure the quality of hash codes. As shown in Figure 1 (a), the general paradigm for obtaining hash codes involves extracting data features using a deep neural network and acquiring hash codes through a hash layer [^1] Most deep hashing methods for image retrieval adhere to this paradigm, while some works (Shen et al., 2017; Jiang and Li, 2018; Chen et al., 2019; Wu et al., 2023) optimize the hash code in the database independently. Our work focuses on the former.. The hash layer usually consists of a single-layer perceptron to map the data features to a length equivalent to that of the hash codes and an operation (e.g., the signum function) for acquiring the binary hash codes.
Figure 1. (a) The general paradigm for deep supervised hashing. We require training multiple models if we need hash codes with varying lengths for effectiveness and efficiency trade-off selection. (b) The Nested Hash Layer can concurrently generate hash codes of varying lengths in a nested fashion while only training once.
While deep supervised hashing methods have achieved notable success, they solely focus on improving the quality of hash codes with a specific code length. As shown in Figure 1 (a), each model corresponds to hash codes of a specific length. This paradigm leads to two problems. On the one hand, a trade-off exists between the efficiency and effectiveness of hash codes. Within a reasonable range, shorter hash codes improve efficiency but compromise effectiveness (Sun et al., 2023). Conversely, longer hash codes generally yield superior performance at the cost of heightened storage and computational overhead. Therefore, we must train multiple deep hashing models corresponding to different code lengths to select the most suitable code length for the current task. This will escalate training time and computational resources expenditure (Wu et al., 2022). On the other hand, as the current deep hashing models can only generate hash codes of a single length, the potential relationships among hash codes of varying lengths are thus disregarded.
In this work, we tackle the above problems by introducing a plug-and-play module, the Nested Hash Layer (NHL), to replace the original hash layer in deep hashing models. First, we observe that the same backbone is employed in deep hashing to extract data features regardless of the chosen hash code length. Furthermore, longer hash codes can be viewed as supplementary descriptions of shorter ones. For example, within an 8-bit hash code, the first 4 bits can be regarded as a 4-bit hash code, while the subsequent 4 bits can be seen as complementary descriptions. Based on the two observations, the basic structure of the NHL has been proposed as shown in Figure 1 (b). This structure aims to concurrently generate hash codes of varying lengths in a nested fashion, which allows for obtaining hash codes of varying lengths from a single training session. Second, the basic NHL combines objectives of varying code lengths, but it cannot ensure different objectives always mutually reinforce each other. Moreover, the objectives associated with shorter hash codes become more critical as they form a part of the longer hash codes when applying NHL. This complexity underscores the need for careful consideration. To address this problem, we apply a heuristic adaptive weights strategy to the basic NHL. Specifically, we first propose the concept of domination gradient for each nested parameter in NHL. The domination gradient serves as a direction for the optimization of the corresponding parameter, prioritizing the short hash codes. Then, we monitor gradients on parameters and dynamically adjust the weights of objectives to ensure that the final optimization direction does not conflict with the domination gradient. Third, in contrast to the traditional hash layer, NHL can simultaneously acquire multiple lengths of hash codes. Recognizing that the relationships inherent in long hash codes can guide those of short hash codes, we propose a long-short cascade self-distillation to improve the hash codes’ quality further. In summary, we make the following contributions:
  • We propose a plug-and-play module called Nested Hash Layer (NHL), which can concurrently generate hash codes of varying code lengths solely by replacing the hash layer of deep supervised hashing models.
  • Based on the Nested Hash Layer (NHL) architecture, we address the gradient conflicts problem by incorporating an adaptive weights strategy for learning objectives and introduce a long-short cascade self-distillation to enhance the quality of the hash codes.
  • We perform comprehensive experiments about effectiveness and efficiency with various deep hashing models. The results indicate that the NHL can achieve an overall training speed increase of around 5-8 times while ensuring effective retrieval outcomes.
2. Related work
2.1. Deep Supervised Hashing
@@
The prevailing deep supervised hashing can be roughly divided into pair-wise methods (Liu et al., 2016; Zhu et al., 2016, 2017; Cao et al., 2018; Li et al., 2020; Zheng et al., 2020), ranking-based methods (Wang et al., 2017; Liu et al., 2018; Cakir et al., 2019), and proxy-based methods (Yuan et al., 2020; Fan et al., 2020; Hoe et al., 2021; Wang et al., 2022, 2023). The objective of pair-wise methods is to ensure similar pairs have similar hash codes while dissimilar pairs have dissimilar hash codes. Ranking-based methods adopt ranking-based similarity-preserving loss terms. For instance, triplet loss (Wang et al., 2017; Liu et al., 2018) and list-wise loss (Cakir et al., 2019) are commonly used to maintain data ordering. Proxy-based methods, also known as center-based methods, have emerged as a widely acclaimed approach recently. These methods first generate each category’s proxies (or hash centers). Then, they force hash codes outputted from the network to approach corresponding proxies/centers. For example, CSQ (Yuan et al., 2020) generates hash centers using the Hadamard matrix or Bernoulli sampling. Based on this foundation, SHCIR (Wang et al., 2022) and MDSH (Wang et al., 2023) further incorporate semantic information of classes and impose minimum distance constraints on the hash centers, respectively.
Current deep supervised hashing methods have achieved significant success, but they only consider a model with a specific code length. This limitation leads to slow training in practical applications due to the effectiveness and efficiency trade-off with code lengths. Some works aim to create new codes through code expansion (Mandal et al., 2019; Wu et al., 2022, 2024) or compression (Zhao et al., 2020). However, their focus lies in training a mapping post-model training to transform old hash codes into new hash codes. One potentially relevant approach is MAH (Luo et al., 2020). However, it is limited to learning hash codes of only three different lengths and focuses on enhancing the accuracy of short hash codes. In contrast, our approach directly generates hash codes of arbitrary lengths during training and focuses on enhancing the overall training speed, thus presenting a fundamental distinction.
2.2. Multi-task Learning
The NHL can be partially viewed as a multi-task learning framework (Lee and Seok, 2023), which represents a learning paradigm where multiple distinct yet correlated tasks are simultaneously trained with a shared model. The fundamental role of our proposed NHL is to share a hash model among learning objectives of varying code lengths.
@@ 2.2. Multi-task Learning
A popular line in multi-task learning is architecture design methods, such as hard parameter sharing methods (Kokkinos, 2017; Bragman et al., 2019) and soft parameter sharing methods (Ruder et al., 2019; Gao et al., 2020; Liu et al., 2019). To cater to a wide range of deep hashing models, NHL only makes simple modifications to the hash layer. The basic structure of NHL is partially inspired by RML (Kusupati et al., 2022), which aims to generate representations of various lengths for diverse downstream tasks. However, RML ignores the impact of gradient conflicts and the relationships between representations of different lengths.
@@
To address gradient or task conflicts, some methods re-weight the task losses based on specific criteria such as uncertainty (Kendall et al., 2018), gradient norm (Chen et al., 2018), or difficulty (Guo et al., 2018). Other methods leverage gradient information to modify the gradient on the parameter update procedure. For example, PCGrad (Yu et al., 2020) projects each task gradient to the normal plane of other task gradients before combining them to form the final update vector. GradDrop (Chen et al., 2020) randomly drops out task gradients based on how much they conflict. CAGrad (Liu et al., 2021) ensures convergence to a minimum of the average loss across tasks by gradient manipulation. Nevertheless, these multi-task learning methods assume the importance of different objectives is equivalent. In NHL, the weights of objectives are different because the short hash codes appear to hold greater significance. Resolving this problem remains a further exploration.
3. METHODOLOGY
3.1. Problem Definition
@@
3.2. Hash Code Generation
@@
The process of acquiring hash codes of most deep supervised hashing methods is divided into two parts. First, a deep neural network is employed to extract the feature $v=F(x)\in\mathbb{R}^{l}$ given the data $x_{i}\in X$, where $l$ is the dimension of $v$. Then, a hash layer $H(\cdot)$ is utilized to derive the hash code $h=H(v)$. In most cases, the hash layer consists of a single-layer perceptron to map the data features to a length equivalent to that of the hash code, and an operation $\phi$ for acquiring the hash codes. The whole process to get the hash code $h$ can be formulated as follows:
(1) $h=f(x)=\phi(WF(x)+c),$
@@
where $W\in R^{b\times l}$ and $c\in R^{b}$ are the parameters in the single-layer perception to be learned. Current deep hashing methods usually predefine a code length $b_{k}$ and then train a hash model $f_{k}$ accordingly. However, in practice, the selection of an appropriate code length depends on the specific task at hand, which means we need to train multiple deep hashing models $\{f_{k}\}_{k=1}^{m}$ for different code lengths and select the most suitable one. Such an approach will increase both the training time and computational resources required. To solve this problem, we introduce NHL to replace the original hash layer in deep hashing models. In the following section, we omit the bias $c$ and the operation $\phi$ for conciseness.
3.3. Basic Structure of Nested Hash Layer
@@
Although the predefined code lengths are varying, the same backbone is employed for a specific deep hashing model. Inspired by this observation, we propose the basic structure of NHL to help deep hashing models generate hash codes with varying code lengths in one training procedure.
@@
Figure 2. (a): The basic Nested Hash Layer (NHL) can generate $m$ (here, $m=3$) hash codes with varying lengths in one training procedure. (b) The illustration of adaptive weight strategy. Taking the $W^{(1)}$ as an example. (c) The Long-short Cascade Self-distillation transfer relationship from long hash codes to short hash codes.
@@
As shown in Figure 2 (a), NHL uses a nested parameter $\{W^{(k)}\}_{k=1}^{m}$ to achieve this goal without adding additional parameters to the neural network. The parameter $W^{(k)}=W_{[1:b_{k}]}^{(m)}\in\mathbb{R}^{l\times b_{k}}$ is in a nested fashion, which means $W^{(k)}\subset W^{(k+1)}$. It uses the first $b_{k}$ vectors of the parameter $W^{(m)}\in\mathbb{R}^{l\times b_{m}}$. We can obtain the hash codes with varying lengths $\{h^{(k)}\}_{k=1}^{m}$ through $h^{(k)}=\phi(W^{(k)}v^{T})$. Then, we aim to minimize the following objective.
(2) $\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m}\mathcal{L}_{k}(h_{i}^{(k)},y_{i};\theta_{F},W^{(k)})$
@@
where $\theta_{F}$ is the parameter of backbone, and $\mathcal{L}_{k}$ is the objective of a specific deep hashing model for code length $b_{k}$. In most deep hashing models, $\mathcal{L}_{k}$ can be a combination of multiple objectives, such as the central similarity loss and quantization loss. As it simply involves adding the original objective of the deep hashing model, it does not alter the original optimization method. By minimizing Eq.(2), we force hash codes with varying lengths to ensure their performance.
3.4. Adaptive Weights Strategy
@@
Although basic NHL can generate hash codes with varying lengths, we are unable to predict whether the gradients for different objectives $\{\mathcal{L}_{k}\}_{k=1}^{m}$ are mutually beneficial or detrimental. For example, in the left part of Figure 2 (b), the parameter $W^{(1)}$ is updated by three gradients $g_{1}^{(1)}=\frac{\partial\mathcal{L}_{1}}{\partial W^{(1)}}$, $g_{2}^{(1)}=\frac{\partial\mathcal{L}_{2}}{\partial W^{(1)}}$, and $g_{3}^{(1)}=\frac{\partial\mathcal{L}_{3}}{\partial W^{(1)}}$. Due to the impact of $g_{2}^{(1)}$, $g_{3}^{(1)}$, optimizing the parameters tends to proceed in a direction unfavourable to $g_{1}^{(1)}$ because the negative inner product between $g_{1}^{(1)}$ and $g_{2}^{(1)}$, $g_{3}^{(1)}$. However, The quality of the hash code $h^{(1)}$ is determined by objective $\mathcal{L}_{1}$, which updates $W^{(1)}$ using the gradient $g_{1}^{(1)}$ based on the target’s outcomes. Therefore, if the final optimization direction of $W^{(1)}$ diverges from $g_{1}^{(1)}$, it is highly probable that such a deviation will lead to a deterioration in the quality of $h^{(1)}$ because the wrong optimization direction for it.
@@
Some multi-task learning works (Yu et al., 2020; Chen et al., 2020; Liu et al., 2021; Guangyuan et al., 2022) propose modifying the gradient on the parameter update procedure to prevent gradient conflicts. However, there exists a difference between these multi-task learning settings and NHL. Multi-task learning treats diverse learning objectives as equally important, aiming to balance various learning objectives. In NHL, the objectives corresponding to shorter hash codes appear to hold greater significance, as shorter hash codes are shared by a larger number of longer hash codes. To address this problem, we propose a heuristic adaptive weights strategy that adjusts the weight of each objective $\mathcal{L}_{k}$ by monitoring the gradients. Intuitively, shorter hash codes should be granted higher optimization priority. Based on this motivation, we introduce the following definitions.
@@
Assume the gradient of $\mathcal{L}_{i}$ for $W^{(k)}$ is denoted as $g_{i}^{k}=\frac{\partial\mathcal{L}_{i}}{\partial W^{(k)}}$. We define $g_{k}^{(k)}=\frac{\partial\mathcal{L}_{k}}{\partial W^{(k)}}$ is the dominant gradient, and $k=1,2...,m$. For example, $g_{1}^{(1)}$ is the dominant gradient in Figure 2 (b).
@@
Assume the gradient of $\mathcal{L}$ for $W^{(k)}$ is $g^{k}=\frac{\partial\mathcal{L}}{\partial W^{(k)}}$. We define anti-domination for the update of $W^{(k)}$ if the inner product is negative between $g^{(k)}$ and the dominant gradient $g_{k}^{(k)}$, whereas a positive inner product is termed align-domination.
The dominant gradient $g_{k}^{(k)}$ serves as a guiding principle for the optimization of the parameter $W^{(k)}$. Anti-domination and align-domination are thus employed to ascertain whether the update result of $W^{(k)}$ is congruent with or divergent from the dominant gradient $g_{k}^{(k)}$. For example, the left part of Figure 2 (b) shows that the update of $W^{(1)}$ is anti-domination because the negative inner product between the gradient $g^{(1)}$ and $g_{1}^{(1)}$. We conducted an analysis to observe the occurrence of anti-domination as training progressed. Figure 3 depicts the likelihood of anti-domination occurring about parameter $W^{(1)}$ at each epoch. These results reveal the probability of anti-domination steadily rises over time, eventually stabilizing at a level exceeding 90%. This trend signifies a growing prevalence of anti-domination scenarios as the training progresses.
@@
Figure 3. The probability of anti-domination occurring on the parameter $W^{(1)}$ at each epoch. We set the code lengths $m=5$ and use CSQ as the deep hashing models on three datasets. This trend signifies a growing prevalence of anti-domination scenarios as the training progresses.
Our goal is to avert anti-domination for each $W^{(k)}$, for which we propose an adaptive weights strategy to ensure its realization. Specifically, as shown in the right part of Figure 2 (b), the basic idea is to decrease the magnitudes of gradient $g_{2}^{(1)}$ and $g_{3}^{(1)}$, while increasing the magnitudes of gradient $g_{1}^{(1)}$. Then, $g_{1}^{(1)}$ can dominate the optimization direction. To achieve this goal, we tend to dynamically adjust the weight $\alpha_{k}$ for each objective $\mathcal{L}_{k}$, then the objective Eq. (2) becomes follows:
(3) $\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m}\alpha_{k}\mathcal{L}_{k}(h_{i}^{(k)},y_{i};\theta_{F},W^{(k)})$
To determine the value of $\alpha_{k}$, we first present our objective in a formal manner. Similar to (Chen et al., 2018), we don’t consider the full network weights and focus on the parameter in NHL, i.e., the $\{W^{(k)}\}_{k=1}^{m}$. Assume $\theta_{ij}^{(k)}$ is the angle between two gradients $g_{i}^{(k)}$ and $g_{j}^{(k)}$. For any $k$, we must ensure the following linear programming:
(4) $\alpha_{k}\lVert g_{k}^{(k)}\rVert+\sum_{i>k}\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq 0;\ k\leq m,$

### v2

# Nested Hash Layer: A Plug-and-play Module for Multiple-length Hash Code Learning

@@
Abstract
Deep supervised hashing is essential for efficient storage and search in large-scale image retrieval. Traditional deep supervised hashing models generate single-length hash codes, but this creates a trade-off between efficiency and effectiveness for different code lengths. To find the optimal length for a task, multiple models must be trained, increasing time and computation. Furthermore, relationships between hash codes of different lengths are often ignored. To address these issues, we propose the Nested Hash Layer (NHL), a plug-and-play module for deep supervised hashing models. NHL generates hash codes of multiple lengths simultaneously in a nested structure. To resolve optimization conflicts from multiple learning objectives, we introduce a dominance-aware dynamic weighting strategy to adjust gradients. Additionally, we propose a long-short cascade self-distillation method, where long hash codes guide the learning of shorter ones, improving overall code quality. Experiments indicate that the NHL achieves an overall training speed improvement of approximately 5 to 8 times across various deep supervised hashing models and enhances the average performance of these models by about 3.4%.
1 Introduction
@@
With the growing amount of visual data on the Internet, existing databases are becoming vast. To manage this data in large-scale image databases, hashing represents images as binary hash codes for efficient storage and search luo2023survey . Recently, deep supervised hashing has made significant progress by extracting deep features and using supervised signals to enhance hash code quality. As shown in the upper part of Figure 1(a), the traditional approach involves using a deep neural network to extract features and a hash layer to generate hash codes [^1] Most deep hashing methods for image retrieval adhere to this paradigm, while some works shen2017deep ; jiang2018asymmetric ; chen2019deep ; wu2023deep optimize the hash code in the database independently. Our work focuses on the former.. The hash layer typically consists of a single-layer perceptron that maps features to the desired hash code length, followed by a binary operation (e.g., the signum function) to produce the final hash codes.
Figure 1: (a) A schematic comparison between traditional hash layer methods and our proposed NHL. The NHL can generate hash codes of multiple lengths simultaneously in a nested manner. (b) The performance of four deep supervised hashing models on the CIFAR-10 dataset highlights the uncertainty in effectiveness and efficiency under different code lengths.
However, most deep supervised hashing models focus on generating hash codes of a specific length. This leads to two problems. First, Figure 1(b) shows the performance of four deep supervised hashing models liu2016deep ; wang2017deep ; cao2018deep ; wang2022image on the CIFAR-10 dataset at different code lengths. There is a clear trade-off between efficiency and effectiveness: shorter hash codes improve efficiency but reduce effectiveness, while longer hash codes enhance performance but increase storage and computational costs sun2023stepwise . Furthermore, this trade-off is not entirely inversely proportional, as seen with the DSH model, which suffers a notable performance drop at 128 bits, a phenomenon known as the dimension curve of hash code. As a result, multiple models must be trained for different code lengths to find the best fit for a specific task, greatly increasing training time and resource use wu2022efficient . Second, since these models only produce single-length hash codes, they overlook the potential relationships between hash codes of different lengths. This raises a question: is it possible to train a single hashing model capable of producing hash codes of multiple lengths?
Code expansion and compression-oriented deep hashing methods address the variation in hash code lengths after training. They generate new codes either through code expansion mandal2019growbit ; wu2022efficient ; wu2024pairwise or compression zhao2020asymmetric . However, their primary focus is on learning a mapping after model training to convert existing hash codes into new ones. MAH luo2020collaborative and SDMLH nie2022supervised involve the generation of hash codes with multiple lengths, but they rely on specifically designed models that cannot be generalized to other deep hashing models. Thus, designing a method capable of being widely applied to deep hashing models for generating hash codes of multiple lengths remains an unexplored area.
In this work, we propose the Nested Hash Layer (NHL), a plug-and-play module that replaces the traditional hash layer in deep supervised hashing models to generate hash codes of multiple lengths. First, we observe that deep supervised hashing models use the same backbone to extract features, regardless of code length. Additionally, longer hash codes can be seen as extensions of shorter ones. Based on these insights, the NHL is designed as shown in the lower part of Figure 1(a) to generate hash codes of different lengths in a nested manner, enabling multiple-length code generation in one model. Second, while the NHL combines objectives for multiple code lengths, conflicts may arise, as shorter hash codes are integral to longer ones. Thus, we introduce a Dominance-Aware Dynamic Weighting strategy. We define a "domination gradient" for each nested parameter, prioritizing the optimization of shorter hash codes. By monitoring parameter gradients, we dynamically adjust objective weights to align with the domination gradient and avoid conflicts. Third, unlike the traditional hash layer, NHL generates multiple-length hash codes. To further enhance code quality, we propose a Long-short Cascade Self-distillation method, leveraging the relationships in long hash codes to guide and improve shorter ones.
As evidenced by extensive experiments and analysis, NHL offers the following key advantages: (1) NHL achieves an overall training speed improvement of approximately 5 to 8 times across various deep supervised hashing models. (2) While ensuring faster training, NHL enhances the average performance of these models by about 3.4% and remarkably alleviates the dimensionality curse of hash codes. (3) NHL demonstrates exceptional flexibility in adapting to scenarios with multiple hash code length settings and different backbones.
2 Related work
2.1 Deep Supervised Hashing
@@
Our work focuses on deep supervised hashing models, which can be roughly divided into pair-wise methods, ranking-based methods, and proxy-based methods. The objective of pair-wise methods liu2016deep ; zhu2016deep ; zhu2017locality ; cao2018deep ; li2020push ; zheng2020deep is to ensure similar pairs have similar hash codes while dissimilar pairs have dissimilar hash codes. Ranking-based methods adopt ranking-based similarity-preserving loss terms. For instance, triplet loss wang2017deep ; liu2018deep and list-wise loss cakir2019hashing are commonly used to maintain data ordering. Proxy-based methods yuan2020central ; fan2020deep ; hoe2021one ; wang2022image ; wang2023deep , also known as center-based methods, have emerged as a widely acclaimed approach recently. These methods first generate each category’s proxies (or hash centers). Then, they force hash codes outputted from the network to approach corresponding proxies (or hash centers). Most of the current deep supervised hashing models only account for a single model with a specific code length. This limitation leads to slow training in practical applications due to the need to train multiple models with different hash code lengths. MAH luo2020collaborative and SDMLH nie2022supervised attempt to solve this problem, but they rely on specifically designed models that cannot be generalized to other deep hashing models.
2.2 Multi-task Learning
@@
The NHL can be seen as a multi-task learning framework lee2023multi , where multiple related tasks are trained simultaneously using a shared model. Its primary function is to enable a single hash model to serve multiple learning objectives for different code lengths. A key aspect of multi-task learning is architecture design, including hard parameter sharing methods kokkinos2017ubernet ; bragman2019stochastic and soft parameter sharing methods ruder2019latent ; gao2020mtl ; liu2019end . NHL only makes simple adjustments to the hash layer to accommodate various deep hashing models. MRL kusupati2022matryoshka partly inspired its basic structure, which generates representations of different lengths for multiple tasks. However, MRL does not address gradient conflicts or the relationships between representations of different lengths.
@@
To address gradient or task conflicts, some methods re-weight the task losses based on specific criteria such as uncertainty kendall2018multi , gradient norm chen2018gradnorm , or difficulty guo2018dynamic . Other methods leverage gradient information to modify the gradient on the parameter update procedure yu2020gradient ; chen2020just ; liu2021conflict ; javaloy2022rotograd . Nevertheless, these multi-task learning methods assume the importance of different objectives is equivalent. In NHL, the weights of objectives are different because the short hash codes appear to hold greater significance. Resolving this problem remains a further exploration.
3 METHODOLOGY
3.1 Problem Definition
@@
3.2 Hash Code Generation
@@
The process of acquiring hash codes of most deep supervised hashing methods is divided into two parts. First, a deep neural network is employed to extract the feature $v=\mathcal{F}(x)\in\mathbb{R}^{l}$ given the data $x_{i}\in X$, where $l$ is the dimension of $v$. Then, a hash layer is utilized to derive the hash code $h$. In most cases, the hash layer consists of a single-layer perceptron to map the data features to a length equivalent to that of the hash code, and an operation $\phi$ for acquiring the hash codes. The whole process to get the hash code $h$ can be formulated as follows:
$h=f(x)=\phi(\mathcal{W}\mathcal{F}(x)+c),$ (1)
@@
where $\mathcal{W}\in R^{b\times l}$ and $c\in R^{b}$ are the parameters in the single-layer perception to be learned. Current deep hashing methods usually predefine a code length $b_{k}$ and then train a hash model $f_{k}$ accordingly. However, in practice, the selection of an appropriate code length depends on the specific task at hand, which means we need to train multiple deep hashing models $\{f_{k}\}_{k=1}^{m}$ for different code lengths and select the most suitable one. Such an approach will increase both the training time and computational resources required. To solve this problem, we introduce NHL to replace the original hash layer in deep hashing models. In the following section, we omit the bias $c$ and the operation $\phi$ for conciseness.
3.3 Basic Structure of Nested Hash Layer
@@
Although the predefined code lengths differ, the same backbone is employed for a specific deep hashing model. Inspired by this observation, we propose the basic structure of NHL to help deep hashing models generate hash codes with different code lengths in one training procedure.
@@
Figure 2: (a): The Nested Hash Layer (NHL) can generate $m$ (here, $m=3$) hash codes with varying lengths in one training procedure. (b) The illustration of the Dominance-Aware Dynamic Weighting strategy. Taking the $\mathcal{W}^{(1)}$ as an example. (c) The Long-short Cascade Self-distillation transfer relationship from long hash codes to short hash codes.
@@
As shown in Figure 2a, NHL uses a nested parameter $\{\mathcal{W}^{(k)}\}_{k=1}^{m}$ to achieve this goal without adding additional parameters to the neural network. The parameter $\mathcal{W}^{(k)}=\mathcal{W}_{[1:b_{k}]}^{(m)}\in\mathbb{R}^{l\times b_{k}}$ is in a nested structure, which means $\mathcal{W}^{(k)}\subset\mathcal{W}^{(k+1)}$. It uses the first $b_{k}$ vectors of the parameter $\mathcal{W}^{(m)}\in\mathbb{R}^{l\times b_{m}}$. We can obtain the hash codes with different lengths $\{h^{(k)}\}_{k=1}^{m}$ through $h^{(k)}=\phi(\mathcal{W}^{(k)}v^{T})$. Then, we aim to minimize the following objective.
$\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m}\mathcal{L}_{k}(h_{i}^{(k)},y_{i};\theta_{F},\mathcal{W}^{(k)}),$ (2)
@@
where $\theta_{F}$ is the parameter of backbone, and $\mathcal{L}_{k}$ is the objective of a specific deep hashing model for code length $b_{k}$. In most deep hashing models, $\mathcal{L}_{k}$ can be a combination of multiple objectives, such as the central similarity loss and quantization loss. As it simply involves adding the original objective of the deep hashing model, it does not alter the original optimization method. By minimizing Eq.(2), we force hash codes with different lengths to ensure their performance.
3.4 Dominance-Aware Dynamic Weighting strategy
@@
Although basic NHL can generate hash codes with different lengths, we are unable to predict whether the gradients for different objectives $\{\mathcal{L}_{k}\}_{k=1}^{m}$ are mutually beneficial or detrimental. For example, in the left part of Figure 2b, the parameter $\mathcal{W}^{(1)}$ is updated by three gradients $g_{1}^{(1)}=\frac{\partial\mathcal{L}_{1}}{\partial\mathcal{W}^{(1)}}$, $g_{2}^{(1)}=\frac{\partial\mathcal{L}_{2}}{\partial\mathcal{W}^{(1)}}$, and $g_{3}^{(1)}=\frac{\partial\mathcal{L}_{3}}{\partial\mathcal{W}^{(1)}}$. Due to the impact of $g_{2}^{(1)}$, $g_{3}^{(1)}$, optimizing the parameters tends to proceed in a direction unfavourable to $g_{1}^{(1)}$ because the negative inner product between $g_{1}^{(1)}$ and $g_{2}^{(1)}$, $g_{3}^{(1)}$. However, The quality of the hash code $h^{(1)}$ is determined by objective $\mathcal{L}_{1}$, which updates $\mathcal{W}^{(1)}$ using the gradient $g_{1}^{(1)}$ based on the target’s outcomes. Therefore, if the final optimization direction of $\mathcal{W}^{(1)}$ diverges from $g_{1}^{(1)}$, it is highly probable that such a deviation will lead to a deterioration in the quality of $h^{(1)}$ because the wrong optimization direction for it.
@@
Some multi-task learning works yu2020gradient ; chen2020just ; liu2021conflict ; javaloy2022rotograd ; guangyuan2022recon propose modifying the gradient on the parameter update procedure to prevent gradient conflicts. However, there exists a difference between these multi-task learning settings and NHL. Multi-task learning treats diverse learning objectives as equally important, aiming to balance various learning objectives. In NHL, the objectives corresponding to shorter hash codes appear to hold greater significance, as shorter hash codes are shared by a larger number of longer hash codes. To address this problem, we propose a dominance-aware dynamic weighting strategy to adjust the weight $\alpha_{k}$ of each objective $\mathcal{L}_{k}$ by monitoring the gradients. Then the objective Eq. (2) becomes follows:
$\mathcal{L}=\sum_{i=1}^{N}\sum_{k=1}^{m}\alpha_{k}\mathcal{L}_{k}(h_{i}^{(k)},y_{i};\theta_{F},\mathcal{W}^{(k)}).$ (3)
Since shorter hash codes should be given higher optimization priority, we are motivated to introduce the following definitions.
@@
Assume the gradient of $\mathcal{L}_{i}$ for $\mathcal{W}^{(k)}$ is denoted as $g_{i}^{k}=\frac{\partial\mathcal{L}_{i}}{\partial\mathcal{W}^{(k)}}$. We define $g_{k}^{(k)}=\frac{\partial\mathcal{L}_{k}}{\partial\mathcal{W}^{(k)}}$ is the dominant gradient, and $k=1,2...,m$. For example, $g_{1}^{(1)}$ is the dominant gradient in Figure 2b.
@@
Assume the gradient of $\mathcal{L}$ for $\mathcal{W}^{(k)}$ is $g^{k}=\frac{\partial\mathcal{L}}{\partial\mathcal{W}^{(k)}}$. We define anti-domination for the update of $\mathcal{W}^{(k)}$ if the inner product is negative between $g^{(k)}$ and the dominant gradient $g_{k}^{(k)}$, whereas a positive inner product is termed align-domination.
The dominant gradient $g_{k}^{(k)}$ serves as a guiding principle for the optimization of the parameter $\mathcal{W}^{(k)}$. Anti-domination and align-domination are thus employed to ascertain whether the update result of $\mathcal{W}^{(k)}$ is congruent with or divergent from the dominant gradient $g_{k}^{(k)}$. For example, the left part of Figure 2b shows that the update of $\mathcal{W}^{(1)}$ is anti-domination because the negative inner product between the gradient $g^{(1)}$ and $g_{1}^{(1)}$. We conducted an analysis to observe the occurrence of anti-domination as training progressed. Figure 3 depicts the likelihood of anti-domination occurring about parameter $\mathcal{W}^{(1)}$ at each epoch. These results reveal the probability of anti-domination steadily rises over time, eventually stabilizing at a level exceeding 90%. This trend signifies a growing prevalence of anti-domination scenarios as the training progresses.
@@
Figure 3: The probability of anti-domination occurring on the parameter $\mathcal{W}^{(1)}$ at each epoch. We set the code lengths $m=5$ and use CSQ as the deep hashing models on three datasets. This trend signifies a growing prevalence of anti-domination scenarios as the training progresses.
Our goal is to avert anti-domination for each $\mathcal{W}^{(k)}$. We propose the following proposition:
Proposition 1.